# Quantum LDPC codes for erasure-biased atomic quantum processors

### v1

@@ III.3 Error correction simulations
The comparison between La-cross and surface codes shows that there always exist physical error probabilities below which La-cross codes offer larger error suppression than the surface code [crossing points of solid and dashed lines of the same color in Fig. 2(a)-(c)]. In Ref. [11], La-cross codes display larger error suppression than the surface code at equal number of physical and logical qubits for physical error probabilities $p\lesssim 10^{-3}$. Instead, in this work we find that, for a $R_{e}=0.98$ unbiased erasure fraction, the onset of improvement in the logical failure probability over the surface code is larger and occurs at higher and experimentally more accessible physical error probabilities. For example, for $D>5$ La-cross codes, the logical error probability at $p\approx 10^{-2}$ is already approximately one order of magnitude lower than that of the surface code. This behavior is due to the higher thresholds and to the different dependence of the logical error probability on the code distance compared to pure depolarizing noise. In fact, for pure erasures, the scaling is $P_{L}\propto p^{D}$, which is a faster decrease of the logical error probability compared to depolarizing noise, for which $P_{L}\propto p^{\lfloor(D+1)/2\rfloor}$. For $R_{e}<1.0$, Pauli errors are dominant in the asymptotic limit, while erasures are dominant near-threshold. Therefore, we expect the logical error probability to scale faster around threshold and then to slightly “bend” asymptotically to produce the expected $P_{L}\propto p^{\lfloor(D+1)/2\rfloor}$ at $p\ll p_{th}$ [26]. This behavior is visible for the small-distance codes shown in Fig. 2(a)-(c).
@@
Figure 4: (a) Cumulative logical error probability normalized by the number of QEC rounds for Bivariate Bicycle (BB) codes under different fractions of unbiased erasure errors and Pauli errors, $R_{e}=0.00-0.98$. For each case, we study three code instances, namely $[[72,12,6]]$, $[[108,8,10]]$ and $[[144,12,12]]$, represented with different gradients of the same color and marker type. Thresholds (black stars) saturate at $\approx 1.1-1.2\%$ for $R_{e}>0.10$, while the logical error probability at the threshold decreases. (b) Comparison of the threshold behavior under different unbiased erasure fractions for BB codes (red), surface code (black) and $k=2,3,4$ La-cross codes (gray). (c) Comparison of the logical error probability normalized by the number of logical qubits, $K$, at $R_{e}=0.98$ for distance $D=6$ BB (red) and La-cross codes (gray). All $k=2,3,4$ La-cross codes are found to offer slightly lower logical failure probabilities than BB codes, while their encoding rate is lower, i.e. $K^{BB}/N^{BB}=12/72$, while $K^{k=2}/N^{k=2}=4/130$, $K^{k=3}/N^{k=3}=9/225$ and $K^{k=4}/N^{k=4}=16/208$ for $D=6$.
@@
The thresholds and the onset of improvement in the logical error probability for La-cross codes over the surface code get even larger under biased erasure errors. We show the decoding plots for $k=2,3,4$ La-cross codes with $R_{e}=0.98$ of biased erasures (dark lines) against unbiased erasures (shaded lines) in Fig. 3(a)-(c). The thresholds approximately increase from $4.0\%$ to $6.8\%$ for $k=2$ La-cross codes, from $4.2\%$ to $7.0\%$ for $k=3$ La-cross codes and from $4.6\%$ to $7.2\%$ for $k=4$ La-cross codes. These values are therefore again comparable to the $10.1\%$ threshold of the surface code under the same noise model, while the offered logical error protection is significantly larger.
@@
To produce the plots with biased erasure errors, we have optimized the scaling factor of the Belief Propagation decoder routine, $s$, to minimize the output logical error probability, finding optimal values around $s=0.3-0.4$. We also note that, due to the lower accuracy of the decoder at large code sizes, we have chosen to identify the thresholds as the crossing point between the $D=4$ and $D=5$ lines [black stars in Fig. 3(a)-(c)].
@@
We show in Tab. 1 a summary of the threshold values for surface and La-cross codes at $R_{e}=0.00,0.98,1.00$ for both unbiased and biased erasures. For the surface code, we also report in parenthesis the threshold values that we have obtained before normalizing the data by the number of QEC rounds, for comparison with the literature (see Appendix A for a more detailed discussion).
@@
We find that, surprisingly, Bivariate Bicycle codes behave differently from La-cross codes under unbiased erasure errors, as shown in Fig. 4(a), where the decoding plots for several erasure fractions are displayed. For $R_{e}=0.10$, the circuit-level threshold improves from $0.7\%$ ($R_{e}=0$) to $1.0\%$ and the logical error probability at the threshold, $P_{L}(p_{th})$, increases by a factor of approximately ×$\times 1.5$. This behavior is similar to the one displayed by surface and La-cross codes with increasing unbiased erasure fractions. In contrast, for larger erasure fractions ($R_{e}=0.50,0.75,0.90,0.95$ in the figure) the threshold does not increase further, but rather seems to saturate around $p_{th}\approx 1.1\%$ [see also Fig. 4(b), red line for Bivariate Bicycle codes] and the logical error probability at the threshold decreases. When $R_{e}=0.98$, the threshold is $p_{th}\approx 1.2\%$, corresponding to an approximate ×$\times 1.7$ improvement compared to pure depolarizing noise, and the logical error probability there is three orders of magnitude lower than that at $R_{e}=0.10$, namely $P_{L}(p_{th})\approx 10^{-4}$.
@@
These results show that different quantum LDPC code families can benefit differently from large erasure fractions. In particular, Bivariate Bicycle codes still do take advantage of large fractions of erasure errors, like surface and La-cross codes, but this advantage manifests itself only as a lower logical error probability, without any significant threshold improvement. This behavior may be relevant for near-term realizations of erasure conversion with quantum LDPC codes in present noisy quantum devices, where high error correction thresholds are highly desirable for QEC scale-up. However, we note that the value of $P_{L}(p_{th})\approx 10^{-4}$ offered by Bivariate Bicycle codes under $R_{e}=0.98$ of unbiased erasures is far below break-even (i.e., $P_{L}(p)=p$) and therefore potentially still interesting for above-threshold near-term QEC experiments. We also note that, despite the different threshold behavior, the logical error probability offered by Bivariate Bicycle codes is comparable to that of La-cross codes under large erasure fractions. In Fig. 4(c), we compare the logical error protection offered by the $D=6$ Bivariate Bicycle code at $R_{e}=0.98$ of unbiased erasures with that offered by $D=6$ $k=2,3,4$ equal distance La-cross codes under the same noise model. The logical error probability of Bivariate Bicycle codes normalized by the number of logical qubits (red line) is slightly higher than that of all La-cross codes (gray lines). That is because below threshold the QEC performance improves with the system size and La-cross codes have a larger number of physical qubits at equal distance and number of logical qubits. In fact, the encoding rate, $K/N$, offered by Bivariate Bicycle codes is larger, although still asymptotically vanishing, $K/N\xrightarrow{N\rightarrow\infty}0$. We conclude that the error correction performance of the two codes is comparable in the range of physical error probabilities $p\lesssim 1\%$.
@@
The reason for not observing any significant increase in the threshold for Bivariate Bicycle codes is likely due to the fact that, under pure depolarizing noise, these codes already show high values of $D\times P_{L}(p_{th})$ (the $D\times$ factor standing for before round normalization), which gets $D\times P_{L}(p_{th})\approx 1$ at $R_{e}=0.10$ and therefore cannot increase further. However, a noise model with a large erasure fraction is significantly easier to decode than pure depolarizing noise and so the codes are expected to offer larger error suppression, i.e. lower logical error probability, as we observe in this work.
@@
Decoding plots of Bivariate Bicycle codes have been obtained by optimizing over the Belief Propagation scaling factor to find the lowest logical error probabilities, finding optimal values for $s=0.3-0.4$.
@@ IV.1 Implementation schemes for quantum LDPC codes with neutral atoms
In the implementation via qubit shuttling, atoms are trapped in movable tweezer traps generated by a crossed 2D acousto-optic deflector and then translated in parallel to execute CZ gates between neighboring data-ancilla atom pairs [10]. This allows for any-to-any connectivity and nearest-neighbor gate execution at the price of long QEC cycle times, on the order of tens of milliseconds, mostly dominated by atom rearrangement time [10].
The static implementation paradigm requires engineering long-range interactions using multiple lasers to enable CZ gates between distant qubits. Transitions to Rydberg-excited energy levels with different principal quantum numbers are required to execute gates for different interatomic distances [11]. This scheme is estimated to be as much as one order of magnitude faster than atom shuttling, although the CZ gate fidelity decreases quasi-linearly with the gate range [11]. The protocol can be further sped up by loading atoms in a folded array configuration to effectively reduce the gate extents. As a consequence, gate infidelities and durations decrease. Array folding has been originally proposed in Ref. [13] for Bivariate Bicycle codes, and it also applies to La-cross codes with open boundary conditions. La-cross codes with periodic boundaries and the standard Shor syndrome extraction scheme would instead suffer from hook errors, as demonstrated in Ref. [11].
@@ IV.2 Equal erasure fraction and range-independent error strength
The assumption of equal erasure fraction $R_{e}$ for all gates can be motivated for both the dynamic and static realization schemes. In the implementation via qubit shuttling, all CZ gates are executed nearest-neighbor. Therefore, if qubit shuttling is supplied with erasure conversion, all gate errors can be modeled as in Ref. [26] and up to $98\%$ of them can be converted into erasures. In the static implementation, it is still in principle possible to remove all the Rydberg population after the gate, either by waiting a longer time (due to the longer Rydberg lifetime of higher-excited states) or via autoionization as discussed in Ref. [26] for short-range gates, once again converting up to $98\%$ of gate errors into erasures.
@@ IV.3 Erasure conversion with quantum LDPC codes
The theoretical erasure conversion scheme requires interleaving ground state measurements after any gate layer of the syndrome extraction circuit. Along with the erased qubit location, that guarantees that also the specific gate that failed is heralded, thus maximizing the noise information and dramatically simplifying the decoding problem. However, integrating erasure conversion into the compilation of a full QEC code – where large numbers of gates have to be executed to correct for errors – is hard. This is due to the fact that quantum LDPC codes typically suffer from a lower degree of parallelism compared to, for example, the surface code. That is detrimental for erasure conversion, as it would require, e.g. massive ground state fluorescence imaging and atom shelving. As a result, the total time budget and physical error probability would increase.
@@ V Conclusions
These results show how the error correction performance of new quantum LDPC code families can benefit from large erasure fractions, as originally demonstrated for the surface code, setting the stage for near-term experimental demonstrations of high-threshold and low-overhead quantum LDPC code memories. Moreover, these results prove that different code families can benefit differently from large erasure fractions and that the main feature of erasure conversion is the drastic diminishment of the logical error probability due to erasures being more amenable to decode than Pauli errors, rather than an increase of the circuit-level threshold, as mostly believed. This pushes the need for identifying erasure-specific QEC resources, not only at the physical level (e.g., fast or robust quantum gates [41, 59, 60]), but also at the logical level, by designing QEC codes that offer optimal performance under erasure errors, trading between overhead reduction, high threshold, good error correction capabilities, and amenable long-range connectivity for near-term implementations.
Although in this work we have mostly focused on erasure conversion to mitigate Rydberg decays in Alkaline-earth(-like) atom qubits, these results may be generalized to other quantum platforms for which erasure conversion protocols have also been demonstrated, such as trapped ions [33] and superconducting qubits [34], and, partially, to other error mechanisms that can be similarly converted into erasures, such as atom loss [22, 29, 31, 32] or mixtures of atom loss and Rydberg decay [61].
@@
Appendix A: Surface code thresholds

### v2

@@ III.3 Error correction simulations
The comparison between La-cross and surface codes shows that there always exist physical error probabilities below which La-cross codes offer larger error suppression than the surface code [crossing points of solid and dashed lines of the same color in Fig. 2(a)-(c)]. In Ref. [11], La-cross codes display larger error suppression than the surface code at equal number of physical and logical qubits for physical error probabilities $p\lesssim 10^{-3}$. Instead, in this work we find that, for a $R_{e}=0.98$ unbiased erasure fraction, the onset of improvement in the logical failure probability over the surface code is larger and occurs at higher and experimentally more accessible physical error probabilities. For example, for $D>5$ La-cross codes, the logical error probability at $p\approx 10^{-2}$ is already approximately one order of magnitude lower than that of the surface code. This behavior is due to the higher thresholds and to the different dependence of the logical error probability on the code distance compared to pure depolarizing noise. In fact, for pure erasures, the scaling is $P_{L}\propto p^{D}$, which is a faster decrease of the logical error probability compared to depolarizing noise, for which $P_{L}\propto p^{\lfloor(D+1)/2\rfloor}$. For $R_{e}<1.0$, Pauli errors are dominant in the asymptotic limit, while erasures are dominant near-threshold. Therefore, we expect the logical error probability to scale faster around threshold, while asymptotically $P_{L}\propto p^{\lfloor(D+1)/2\rfloor}$ at $p\ll p_{th}$ [26] [for the bending visible in Fig. 2(a)-(c), see Appendix A].
@@
Figure 4: (a) Cumulative logical error probability normalized by the number of QEC rounds for Bivariate Bicycle ($B$) codes under different fractions of unbiased erasure errors and Pauli errors, $R_{e}=0.00-0.98$. For each case, we study three code instances, namely $[[72,12,6]]$, $[[108,8,10]]$ and $[[144,12,12]]$, represented with different gradients of the same color and marker type. Thresholds (black stars) saturate at $\approx 1.1-1.2\%$ for $R_{e}>0.10$, while the logical error probability at the threshold decreases. (b) Logical error probability at $R_{e}=0.98$ of unbiased erasures for Bivariate Bicycle (red) and La-cross codes (gray) at similar encoding rates, $R=K/N$. Codes considered are: $[[108,8,10]]$ ($R\approx 0.074$) against $[[52,4,4]]$ ($R\approx 0.077$), $[[149,9,5]]$ ($R\approx 0.060$), and $[[208,16,6]]$ ($R\approx 0.077$). In this case, $B$ codes are expected to quickly outperform La-cross codes below threshold due to the larger code distance at fixed qubit overhead ($D_{BB}=10$ against $D_{la-cross}=4,5,6$ in the shown example). (c) Logical error probability normalized by the number of logical qubits, $K$, at $R_{e}=0.98$ of unbiased erasures for $B$ (red) and La-cross codes (gray) at equal distance, $D=6$. La-cross codes slightly outperform $B$ codes due to the larger qubit overhead. In fact, La-cross encoding rates are lower than $B$ ones, i.e. $R_{BB}\approx 0.167$, while $R_{k=2}\approx 0.031$, $R_{k=3}=0.04$ and $R_{k=4}\approx 0.077$ for $D=6$.
@@
The thresholds and the onset of improvement in the logical error probability for La-cross codes over the surface code get even larger under biased erasure errors. We show the decoding plots for $k=2,3,4$ La-cross codes with $R_{e}=0.98$ of biased erasures (colored lines) against unbiased erasures (pale lines), and surface codes $R_{e}=0.98$ of biased erasures at equal number of physical and logical qubits in Fig. 3(a)-(c). The thresholds approximately increase from $4.0\%$ to $6.8\%$ for $k=2$ La-cross codes, from $4.2\%$ to $7.0\%$ for $k=3$ La-cross codes and from $4.6\%$ to $7.2\%$ for $k=4$ La-cross codes. These values are therefore again comparable to the $10.1\%$ threshold of the surface code under the same noise model, while the offered logical error protection is significantly larger. Consistently with the observations above for the unbiased erasure noise mode, also with biased erasures, La-cross codes outperform surface codes with equal number of physical and logical qubits for sufficiently small error probabilities, i.e. approximately below $p\lesssim 10^{-2}$ for the code distances here considered.
@@
To produce the plots with biased erasure errors, we have optimized the scaling factor of the Belief Propagation decoder routine, $s$, to minimize the output logical error probability, finding optimal values around s=0.30.4s=0.3-0.4. We also note that, due to the lower accuracy of the decoder at large code sizes, we have chosen to identify as an upper bound for the thresholds the crossing point between the $D=4$ and $D=5$ lines [black stars in Fig. 3(a)-(c)].
We observe that, despite the choice of considering Clifford-deformed La-cross codes, most of the gains from the unbiased and biased erasure noise models are found to be independent of the code deformation (we show the plots in Fig. 6(b),(c) in Appendix A). In fact, for both CSS and non-CSS La-cross codes, we have not observed any difference in logical performance when enforcing the same noise model and the same decoding strategy via BP+OSD decoder. We note that this is consistent with our modeling of unbiased erasures as depolarizing errors with known location. On the other hand, although biased erasures are modeled as $Z$-biased errors, the use of native non-bias preserving gates effectively unbias the noise channel. We note that this behavior of La-cross codes is consistent with that of the surface code shown in Ref. [27], where a comparable logical performance is observed for the XZZX and the standard surface code with $R_{e}=0.98$ of biased erasure errors.
@@
We show in Tab. 1 a summary of the threshold values for surface and La-cross codes at $R_{e}=0.00,0.98,1.00$ for both unbiased and biased erasures. For the surface code, we also report in parentheses the threshold values that we have obtained before normalizing the data by the number of QEC rounds, for comparison with the literature (see Appendix B for a more detailed discussion).
@@
We find that, surprisingly, Bivariate Bicycle codes behave differently from La-cross codes under unbiased erasure errors, as shown in Fig. 4(a), where the decoding plots for several erasure fractions are displayed. For $R_{e}=0.10$, the circuit-level threshold improves from $0.7\%$ ($R_{e}=0$) to $1.0\%$ and the logical error probability at the threshold, $P_{L}(p_{th})$, increases by a factor of approximately $\times 1.5$. This behavior is similar to the one displayed by surface and La-cross codes with increasing unbiased erasure fractions. In contrast, for larger erasure fractions ($R_{e}=0.50,0.75,0.90,0.95$ in the figure) the threshold does not increase further, but rather seems to saturate around $p_{th}\approx 1.1\%$ and the logical error probability at the threshold decreases. When $R_{e}=0.98$, the threshold is $p_{th}\approx 1.2\%$, corresponding to an approximate $\times 1.7$ improvement compared to pure depolarizing noise, and the logical error probability there is three orders of magnitude lower than that at $R_{e}=0.10$, namely $P_{L}(p_{th})\approx 10^{-4}$.
@@
We now compare Bivariate Bicycle codes against La-cross codes in terms of logical error correction performance in the sub-threshold regime when most errors are erasures, namely $R_{e}=0.98$. In Fig. 4(b), we compare one exemplary instance of Bivariate Bicycle codes, namely the $[[108,8,10]]$ code, against La-cross codes of similar encoding rate, $R=K/N$, namely the $[[52,4,4]]$, $[[149,9,5]]$ and $[[208,16,6]]$ codes. We find that below threshold, the Bivariate Bicycle code quickly outperforms all the La-cross codes considered for $p\lesssim 0.008$, due to the larger code distance responsible for a steeper logical error scaling. In Fig. 4(c), we instead compare the logical error protection offered by the $D=6$ Bivariate Bicycle code, i.e. $[[72,12,6]]$ at $R_{e}=0.98$ of unbiased erasures with that offered by $k=2,3,4$ equal distance ($D=6$) La-cross codes, i.e. [[130,4,6]],[[225,9,6]],[[2018,16,5]][[130,4,6]],\,[[225,9,6]],\,[[2018,16,5]], under the same noise model. The logical error probability of Bivariate Bicycle codes normalized by the number of logical qubits (red line) is slightly higher than that of all La-cross codes (gray lines). That is because below threshold the QEC performance improves with the system size and La-cross codes have a larger number of physical qubits at the same distance. In fact, the encoding rate, $K/N$, offered by Bivariate Bicycle codes is larger, although still asymptotically vanishing, $K/N\xrightarrow{N\rightarrow\infty}0$, hence the overhead is lower. The error correction performance of these codes at equal distance is therefore comparable in the range of physical error probabilities $p\lesssim 1\%$.
In conclusion, Bivariate Bicycle codes offer a lower threshold than La-cross codes when most errors are erasures. At the same time, erasure conversion comes at the cost of large operation and time overheads, which make higher threshold desirable to enable near-term implementations. This suggests that in near-term noisy quantum devices La-cross codes might be a more promising candidate to realize erasure conversion in quantum LDPC codes. However, the comparison in logical error correction performance shows that Bivariate Bicycle codes allow to target comparable or lower logical error probabilities with significantly lower qubit overhead. Therefore, these codes might ultimately be preferred at sufficiently small physical error probabilities, approximately $p\lesssim 8\times 10^{-3}$. Additionally, we observe that the value of $P_{L}(p_{th})\approx 10^{-4}$ offered by Bivariate Bicycle codes under $R_{e}=0.98$ of unbiased erasures is far below break-even (i.e., $P_{L}(p)=p$) and therefore potentially interesting also for above-threshold near-term QEC experiments.
Figure 5: Threshold probabilities for different fractions, $R_{e}$, of unbiased erasure errors of Bivariate Bicycle (red), $k=2,3,4$ La-cross (gray) and surface codes (black). For La-cross codes, despite the low values for depolarizing errors, the threshold rapidly increase with the increasing erasure fraction and for $R_{e}\gtrsim 0.98$ they become comparable to the high surface code thresholds. Instead, for Bivariate Bicycle codes, the threshold saturates at $p_{th}^{BB}\approx 1.1\%$ and does not display any significant improvement for increasing erasure fractions.
@@
Finally, our results show that different quantum LDPC code families can benefit differently from large erasure fractions. In particular, Bivariate Bicycle codes still do take advantage of large fractions of erasure errors, like surface and La-cross codes, but this advantage manifests itself only as a lower logical error probability, without any significant threshold improvement, as summarized in Fig. 5. The reason for not observing any significant increase in the threshold for Bivariate Bicycle codes is likely due to the fact that, under pure depolarizing noise, these codes already show high values of $D\times P_{L}(p_{th})$ (the $D\times$ factor standing for before round normalization), which gets $D\times P_{L}(p_{th})\approx 1$ at $R_{e}=0.10$ (cf. Fig. 4(a)) and therefore cannot increase further. However, a noise model with a large erasure fraction has a lower noise entropy and is easier to decode than pure depolarizing noise. Therefore, the codes are still expected to offer larger error suppression, i.e. lower logical error probability, as we observe in this work.
@@
Decoding plots of Bivariate Bicycle codes have been obtained by optimizing over the Belief Propagation scaling factor to find the lowest logical error probabilities, finding optimal values for s=0.30.4s=0.3-0.4.
@@ IV.1 Implementation schemes for quantum LDPC codes with neutral atoms
In the implementation via qubit shuttling, atoms are trapped in movable tweezer traps generated by a crossed 2D acousto-optic deflector and then translated in parallel to execute CZ gates between neighboring data-ancilla atom pairs [10]. This allows for any-to-any connectivity and nearest-neighbor gate execution at the price of large numbers of rearrangement steps and long QEC cycle times, on the order of tens of milliseconds, mostly dominated by atom rearrangement time [10].
The static implementation paradigm requires engineering long-range interactions using multiple lasers to enable CZ gates between distant qubits. Transitions to Rydberg-excited energy levels with different higher principal quantum numbers are required to execute gates over different interatomic distances [11]. In fact, using higher principal quantum number Rydberg states enables larger blockade radii that are necessary to implement long-range gates. At the same time, larger spatial blockades result in larger Rydberg cross-talk and lower parallelism. Even when accounting for this lower parallelism, this scheme is estimated to be as much as one order of magnitude faster than atom shuttling, although the CZ gate fidelity decreases quasi-linearly with the gate range [11]. The protocol can be further sped up by loading atoms in a folded array configuration to effectively reduce the gate extents. As a consequence, gate infidelities and durations decrease. Array folding has been originally proposed in Ref. [13] for Bivariate Bicycle codes, and it also applies to La-cross codes with open boundary conditions. La-cross codes with periodic boundaries and the standard Shor syndrome extraction scheme would instead suffer from hook errors, as demonstrated in Ref. [11].
@@ IV.2 Equal erasure fraction and range-independent error strength
The assumption of equal erasure fraction $R_{e}$ for all gates can be motivated for both the dynamic and static realization schemes. In the implementation via qubit shuttling, all CZ gates are executed nearest-neighbor. Therefore, if qubit shuttling is supplied with erasure conversion, all gate errors can be modeled as in Ref. [26] and up to $98\%$ of them can be converted into erasures. Instead, in the static implementation, it is still in principle possible to remove all the Rydberg population after the gate, either by waiting a longer time (due to the longer Rydberg lifetime of higher-excited states) or via autoionization as discussed in Ref. [26] for short-range gates, once again converting up to $98\%$ of gate errors into erasures.
@@ IV.3 Erasure conversion with quantum LDPC codes
The theoretical erasure conversion scheme requires interleaving ground state measurements after any gate layer of the syndrome extraction circuit. This guarantees that the specific gate that failed is heralded with the erased qubit location, thus maximizing the noise information and dramatically simplifying the decoding problem. However, integrating erasure conversion into the compilation of a full QEC code – where large numbers of gates have to be executed to correct for errors – is hard. This is due to the fact that quantum LDPC codes typically suffer from a lower degree of parallelism compared to, for example, the surface code. That is detrimental for erasure conversion, as it would require, e.g. massive ground state fluorescence imaging and atom shelving. As a result, the total time budget and physical error probability would increase.
@@ V Conclusions
These results show how the error correction performance of new quantum LDPC code families can benefit from large erasure fractions, as originally demonstrated for the surface code, setting the stage for near-term experimental demonstrations of high-threshold and low-overhead quantum LDPC code memories. Moreover, these results prove that different code families can benefit differently from large erasure fractions and that the main feature of erasure conversion is the drastic diminishment of the logical error probability, also due to erasures being more amenable to decode than Pauli errors, rather than an increase of the circuit-level threshold, as mostly believed. This pushes the need for identifying erasure-specific QEC resources, not only at the physical level (e.g., fast or robust quantum gates [47, 66, 67]), but also at the logical level, by designing QEC codes that offer optimal performance under erasure errors, trading between overhead reduction, high threshold, good error correction capabilities, and amenable long-range connectivity for near-term implementations.
Although in this work we have mostly focused on erasure conversion to mitigate Rydberg decays in Alkaline-earth(-like) atom qubits, these results may be generalized to other quantum platforms for which erasure conversion protocols have also been demonstrated, such as trapped ions [35, 36, 37] and superconducting qubits [38, 39, 40], and, partially, to other error mechanisms that can be similarly converted into erasures, such as atom loss [22, 29, 33, 34] or mixtures of atom loss and Rydberg decay [68].
Figure 6: Supplementary details on QEC numerical simulations. Cumulative logical error probability normalized by the number of QEC rounds for $k=2$ La-cross codes. (a) Decoding plots performing either $D$ (colored thick lines) or 1010 rounds (dark thin lines) of stabilizer measurements. Except for the smallest code shown, with $D=3$, the values of the logical error probabilities are comparable within the error bars. (b) QEC simulations with $R_{e}=0.98$ of unbiased erasure errors for Clifford-deformed non-CSS La-cross codes (colored thick lines) and the standard CSS La-cross codes (dark thin lines) with $D$ rounds of stabilizer measurements. Simulations show no difference in performance for the two codes, consistently with the fact that the noise model is unbiased. (c) QEC simulations with $R_{e}=0.98$ of biased erasure errors for Clifford-deformed non-CSS La-cross codes (colored thick lines) and the standard CSS La-cross codes (dark thin lines) with $D$ rounds of stabilizer measurements. Simulations show no difference in performance for the two codes, suggesting that at circuit-level, non-bias preserving gates completely unbias the error channel. Code instances shown in these plots are: [[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]][[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]]. Error bars on the data are standard deviations associated with the Monte Carlo QEC simulations. BP+OSD decoder was used.
@@
Appendix A: Details on QEC simulations
In this section, we provide additional details to support the QEC numerical simulations shown in the main text.
V.0.1 Stabilizer measurement rounds
In the main text, we have shown simulations for surface, La-cross and Bivariate Bicycle codes by performing $D$ rounds – as many as the code distance – of stabilizer measurements. This guarantees robustness against measurement errors with good approximation [4]. However, the logical error probability is sensitive to the number of rounds [14]. For a specific code instance, namely Clifford-deformed $k=2$-La-cross codes with $98\%$ of unbiased erasure errors, we show in Fig. 6(a) QEC simulations with $10$ (dark thin lines) and $D$ rounds (colored thick lines). Except for the shorter distance code shown ($D=3$), the logical error probabilities are consistent within the error bars, while crucially the threshold remains the same. This shows that the threshold values presented in the main text are independent of the number of rounds used in the simulations.
V.0.2 Noise bias and code bias-deformation
We now discuss the relation between erasure noise models and bias deformation in more detail. To do so, for the standard CSS non-deformed La-cross codes we repeat the same QEC simulations that we have shown in the main text for their Clifford-deformed counterparts. We focus only on the case $k=2$, with a fraction $R_{e}=0.98$ of both unbiased and biased erasure errors. For $k=3,4$ La-cross codes a similar behavior is expected. We show in Fig. 6(b),(c) the resulting plots for unbiased and biased erasures, respectively. In both cases, the simulations with BP+OSD decoder show no difference in logical performance between CSS and non-CSS codes with both noise models. Unbiased erasures are simulated and decoded as unbiased depolarizing errors with known locations, therefore we expect no difference between CSS and non-CSS La-cross codes, as we correctly observe. Biased erasure are instead simulated and decoded as $Z$-biased errors with known locations, and bias-deformed codes are known to perform better under biased code-capacity and phenomenological noise [23]. However, in the present case, simulations are performed at circuit level with native non-bias-preserving gates. The comparable performance of CSS and non-CSS La-cross codes then signals that the noise channel gets effectively unbiased during the computation when native gates are used. We observe that this behavior is not surprising, as it has previously been observed for the XZZX and standard surface codes in Ref. [27].
Since all the simulations shown in this work have been performed with the same decoding strategy for both CSS and non-CSS codes, the above observations incidentally also validate the comparison between non-CSS La-cross codes and CSS Bivariate Bicycle codes with unbiased erasure errors and the same decoder presented in the main text.
V.0.3 Surface code decoding strategies
Finally, we comment on our choice of decoding surface codes with BP+OSD decoder, instead of using faster matching-based decoders. Upon comparing the performances of BP+OSD and pymatching [69] decoders on the surface code, BP+OSD was found to yield comparable logical error probabilities within the error bars and slightly higher thresholds than pymatching, and so the former was preferred in this context. Although the slightly higher accuracy of BP+OSD comes at the price of a larger time overhead, in this work we are interested in assessing the QEC performance of different quantum memories, which can be efficiently decoded offline. In fact, erasure conversion does not require real-time decoding. That is because the fidelity of the measurements that herald the errors is solely limited by the single-atom fluorescence imaging fidelity, which is practically $1$, with good approximation. In Ref. [32], no scattering errors have been observed in the erasure detection process, thus bounding their probability to $10^{-6}$ per imaging time.
We conclude by discussing the bending shown at low physical error probabilities by the decoding curves of small distance codes, e.g., $D=3,4$, (Fig. 23).
We have performed numerical simulations for the surface code in order to compare the decoding performance of BP+OSD and pymatching decoders. We show in Fig. 7 the result of the decoder comparison for the $[[13,1,3]]$ and $[[25,1,4]]$ surface codes. Although BP+OSD slightly overestimates the logical error probability near-threshold (which could be fixed by adjusting the BP scaling factor in that regime), pymatching shows no bending, but performs comparably or slightly worse within the error bars in the deep sub-threshold regime. In addition, the threshold probability obtained by pymatching is slightly lower than that obtained by BP+OSD (see Appendix B below). We conclude that BP+OSD is a valuable decoding strategy for the surface code in the simulated regime of physical error probabilities.
Figure 7: Surface code QEC simulations with BP+OSD and pymatching decoders when $98\%$ of the errors are unbiased erasures. Overall, the two decoders performs comparably within the errors bars, while BP+OSD yields a slightly higher error threshold and lower logical error probability for $D=3$. This shows that BP+OSD slightly outperforms, or performs comparably to, pymatching in the simulated regime of physical error probabilities. The logical error rate is here not normalized by the number of QEC rounds. Error bars on the data are standard deviations associated with the Monte Carlo QEC simulations.
Appendix B: Surface code thresholds